OpenStudy (jenniferkeryakous):

ASB sold a total of 805 advance tickets for the mr. Huskey show. The student tickets costs $3 each, and the adult tickets cost $4 each. If the total revenue was $2740, how many of each type of ticket were sold?

OpenStudy (anonymous):

no. of student tickets=x no. of adult tickets=y 3x+4y=2740 (3x=cost*no of student tickets) x+y=805 Can u solve this now.

OpenStudy (jenniferkeryakous):

How do i solve for it after i make the equation: 3x+4y=2740?

OpenStudy (anonymous):

From 2nd eq. x=805-y Put x value in 1st eq. 3(805-y)+4y=2740 =>2415-3y+4y=2740 =>y=2740-2415 y=325 And x=805-y=805-325=480 Check, no. of std tickets=480 no. of adult tickets=325 Total no of tickets=480+325=805..Done!!

OpenStudy (jenniferkeryakous):

@probhat thank you this really helped